# Polycyclic codes over serial rings and their annihilator CSS construction

### v1

@@ Abstract
In this paper, we investigate the algebraic structure for polycyclic codes over a specific class of serial rings, defined as $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ for $i\in\{1,\ldots,s\}$ is a monic square-free polynomial. We define quasi-$s$-dimensional polycyclic codes and establish an $R$-isomorphism between these codes and polycyclic codes over $\mathscr{R}$. We provide necessary and sufficient conditions for the existence of annihilator self-dual, annihilator self-orthogonal, annihilator linear complementary dual, and annihilator dual-containing polycyclic codes over this class of rings. We also establish the CSS construction for annihilator dual-preserving polycyclic codes over the chain ring $R$ and use this construction to derive quantum codes from polycyclic codes over $\mathscr{R}$.
@@
journal: Cryptography and Communicationsmmfootnotetext: Maryam Bajalan is supported by the Bulgarian Ministry of Education and Science, Scientific Programme ”Enhancing the Research Capacity in Mathematical Sciences (PIKOM)”, No. DO1-67/05.05.2022.vvfootnotetext: Both authors were partially supported by Grant TED2021-130358B-I00 funded by MCIU/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR”
@@ 1 Introduction
Polycyclic codes are ideals in the quotient of the polynomial ring $A[x]/\langle f(x)\rangle$, where $A$ is a ring and $f(x)$ is a polynomial in $A[x]$, see [1, 5, 6, 23]. Cyclic codes and constacyclic codes over $A$ are particular instances when $f(x)=x^{n}-1$ and $f(x)=x-\lambda$, respectively. A ring is called uniserial if all its submodules are linearly ordered by inclusion, and it is called serial if it is a direct sum of uniserial rings. In this paper, we aim to investigate polycyclic codes in the case $A$ is the serial ring $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial for $i=1,2,\ldots,s$.
There is a vast literature of polycyclic codes over some special classes of rings, that we slightly review here. Alkenani et al. [3] studyed constacyclic codes over the ring $\mathbb{F}_{q}[u_{1},u_{2}]/\langle u_{1}^{2}-u_{1},u_{2}^{2}-u_{2},u_{1}u_{2}-u_{2}u_{1}\rangle$. Dinh et al. [10] investigated constacyclic codes over the ring $\mathbb{F}_{p}+u_{1}\mathbb{F}_{p}+\cdots+u_{s}\mathbb{F}_{p}$, subject to the conditions $u_{i}^{2}=u_{i}$, $u_{i}u_{j}=u_{j}u_{i}=0$, and $p$ is an odd prime. Dinh et al. [12] explored the algebraic structure of constacyclic codes over finite semi-simple rings, defined as direct sums of finite fields. Islam et al. [21] focused their study on cyclic codes over the non-chain ring $\mathbb{F}_{q}[u]/\langle u^{e}-1\rangle$. Bag et.al [4] examined skew constacyclic codes over the ring $\mathbb{F}_{q}[u,v]/\langle u^{2}-1,v^{2}-1,uv-vu\rangle$. Ali et al. [2] worked out the case cyclic codes over non-chain rings $\mathbb{F}_{q}[u_{1},u_{2},\ldots,u_{s}]/\langle u^{2}_{i}-(\alpha_{i})^{2},\quad u_{i}u_{j}-u_{j}u_{i}\rangle$, where each $\alpha_{i}$ is a non-zero element of $\mathbb{F}_{q}$. Goyal and Raka [15, 16] explored polyadic cyclic and constacyclic codes over the non-chain ring $\mathbb{F}_{q}[u,v]/\langle f(u),g(v),uv-vu\rangle$, with $f(u)$ and $g(v)$ splitting into distinct linear factors over $\mathbb{F}_{q}$. Bhardwaj et al. [7] investigated constacyclic codes over the finite non-chain ring $\mathbb{F}_{q}[u_{1},u_{2},\ldots,u_{k}]/\langle f_{i}(u_{i}),u_{i}u_{j}-u_{j}u_{i}\rangle$, where each $f_{i}(u_{i})$ is a polynomials that splits into distinct linear factors over $\mathbb{F}_{q}$. Qi [32] focused on polycyclic codes over $\mathbb{F}_{q}+u\mathbb{F}_{q}$, with $u^{2}=0$. Karthick [22] enquired about polycyclic codes over the ring $R=\mathbb{F}_{q}+u\mathbb{F}_{q}+v\mathbb{F}_{q}$, subject to the conditions $u^{2}=\alpha u$, $v^{2}=v$, and $uv=vu=0$, where $\alpha$ represents a unit element of $R$.
@@
In this paper, we investigate polycyclic codes over the general ring $\mathscr{R}$. The outline of the paper is as follows. Section 2 shows the foundational structure of $\mathscr{R}$. In Section 3, we examine how a polycyclic code $\mathscr{C}$ over $\mathscr{R}$ decomposes into a direct sum $\oplus_{i=1}^{k}\mathscr{C}_{i}$, where each $\mathscr{C}_{i}$ constitutes a polycyclic code over a chain ring. Section 4 explores the concept of quasi-2-dimensional polycyclic (Q2DP) codes along with their multidimensional counterparts (QsDP), which extend the ideas of quasi-polycyclic codes and multidimensional cyclic codes [18, 19, 20, 27, 36]. Furthermore, we define the concept of $f(x)$-polycyclic-QsDP codes as a specific subclass of QsDP codes. Then, we prove that when considering polycyclic codes over $\mathscr{R}$ as $R$-submodules, they precisely coincide with $f(x)$-polycyclic-QsDP codes. Polycyclic codes are defined as the invariant $\mathscr{R}$-submodules of $\mathscr{R}^{n}$ under the operator $\tau_{f(x)}$. Their counterpart are sequential codes, which are defined as invariant subspace under the adjoint operator of $\tau_{f(x)}$. It is well-known that if a linear subspace is invariant under one operator, its Euclidean dual is invariant under the adjoint operator. Consequently, a polycyclic code $\mathscr{C}$ and its Euclidean dual $\mathscr{C}^{\perp}$ do not share the same structure. Thus, the first ones are equivalent to ideals in $\mathscr{R}[x]/\langle f(x)\rangle$, defining a polynomial and ideal structure for the second ones is challenging. Note that, if $\mathscr{C}$ is an ideal in $\mathscr{R}[x]/\langle f(x)\rangle$ generated by the polynomial $g(x)$ with the property that $f(x)=g(x)h(x)$, it cannot be assumed that $\mathscr{C}^{\perp}$ is generated by $h(x)$ because $\mathscr{C}^{\perp}$ is not necessarily a polycyclic code. This incorrect interpretation happened in the proof of Theorem 5 in [30]. In Section 5, using multivariable codes over chain rings in [24, 25], we establish a polynomial structure for $\mathscr{C}^{\perp}$ for the special case where $f(x)$ is a polynomial with coefficients in $R$. The annihilator dual is an alternative duality concept ensuring that the dual of polycyclic codes has a polycyclic structure, a feature that fails in Euclidean dual. In Section 6, some properties of the annihilator dual of a polycyclic code over $\mathscr{R}$ are studied. Usually, it is stated in the literature that “a polycyclic code $\langle g(x)\rangle$ is annihilator self-dual if and only if $f(x)=g^{2}(x)$. As explained in Remark 5, that statement relies on the assumption that “if $b$ divides $c$ and $c$ divides $b$, then $b$ equals $c$’, without considering that the divisibility relation for rings is different, namely “if $b$ divides $c$ and $c$ divides $b$, then $\langle b\rangle$ equals $\langle c\rangle$. Thus in Section 6, we revise this property as follows, “a polycyclic code $\langle g(x)\rangle$ is annihilator self-dual if and only if $f(x)=ag^{2}(x)$ for some unit element $a$ in the ring $\mathscr{R}$. To finish this paper, we study some quantum code constructions from this type of codes. Calderbank-Shor-Steane (CSS) construction allows for constructing quantum codes from dual-preserving classical codes, see for example [29, 33]. Existing literature has applied this construction only to codes that preserve Euclidean or Hermitian duality. In this paper, we present construction based on preserving the annihilator duality. In Section 7, we provide an annihilator CSS construction to derive quantum codes from annihilator dual-preserving $f(x)$-polycyclic codes generated by a single element $g(x)$, where $g(x)$ divides $f(x)$.
@@
Let $R$ be a finite commutative chain ring, and let $\gamma\in R$ be a fixed generator of the maximal ideal $\mathfrak{m}$ with nilpotency index $e$. Consider $\mathbb{F}_{q}=R/\mathfrak{m}$ as the residue field, where $q=p^{t}$, $p$ a prime number. Let ¯$\overline{\phantom{x}}:R[y]\to\mathbb{F}_{q}[y]$ represent the natural ring homomorphism, mapping $r\in R$ to $r+\mathfrak{m}$ and the variable $y$ to itself. A polynomial $p(y)=\sum_{i=0}^{n-1}p_{i}y^{i}\in R[y]$ is called square-free (basic irreducible) if $\bar{p}(y)$ is square-free (basic irreducible) in $\mathbb{F}_{q}[y]$. Additionally, it is called regular if it is not a zero divisor in $R[y]$. In [11, Proposition 2.3] it is stated that $p(y)$ is regular if and only if $p_{i}$ is a unit element for some $i$. Hence, monic polynomials are regular. If $g(y)\in R[y]$ is a regular basic irreducible polynomial, then the quotient ring $R[y]/\langle g(y)\rangle$ forms a chain ring, see [11, Lemma 3.1]. It is well-known that the monic square-free polynomial $g(y)$ factors uniquely as a product of monic basic irreducible pairwise coprime polynomials, given by $g(y)=g_{1}(y)g_{2}(y)\cdots g_{r}(y)$[28, Theorem XIII.4]. Hence, by the Chinese Remainder Theorem, the quotient ring $R[y]/\langle g(y)\rangle$ can be decomposed as a direct sum of chain rings
@@
The algebraic properties of the algebra $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, assuming that each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial of degree $m_{i}$, was explored in [24]. Let $H_{i}$ denote the set of roots of $\bar{t}_{i}(x_{i})$ in a suitable extension of $\mathbb{F}_{q}$. For each $\nu\in\mathcal{H}=\prod_{i=1}^{s}H_{i}$, let $C(\nu)$ denote the set of the class of $\nu$ as $\{(\nu^{q^{j}_{1}},\ldots,\nu^{q^{j}_{s}})\mid j\in\mathbb{N}\}$. The set of all these classes, denoted by $\mathcal{C}=\mathcal{C}(t_{1},\ldots,t_{s})$, forms a partition of $\mathcal{H}$. The size of each class is given by $|C(\nu)|=\text{lcm}(d_{1},\ldots,d_{s})=[\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{s}):\mathbb{F}_{q}]$, where $d_{i}$ is the degree of the irreducible polynomial of $\nu_{i}$ over $\mathbb{F}_{q}$. Let $p_{C,i}(x_{i})$ denote the polynomial $\text{Irr}(\nu_{i},\mathbb{F}_{q})$, where $\nu=(\nu_{1},\ldots,\nu_{s})\in\mathcal{H}$ and $C=C(\nu)$ is a class of $\nu$. For each $i=2,\ldots,s$, let $b_{C,i}(x_{i})=\text{Irr}(\nu_{i},\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1}))\in\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1})[x_{i}]$ and $\widetilde{b}_{C,i}(x_{i})=\frac{p_{C,i}(x_{i})}{b_{C,i}(x_{i})}$. The polynomials $b_{C,i}(x_{i})$ and $\widetilde{b}_{C,i}(x_{i})$ are coprime polynomials. One can define the multivariable polynomials $w_{C,i}(x_{1},\ldots,x_{i})$ and $\pi_{C,i}(x_{1},\ldots,x_{i})$ obtained by substituting $\nu_{i}$ by $x_{i}$ in polynomials $b_{C,i}(X_{i})$ and $\widetilde{b}_{C,i}(x_{i})$, respectively and denote the Hensel lifts to $R$ of the polynomials $p_{C,i}$, $w_{C,i}$, and $\pi_{C,i}$ by $q_{C,i}$, $z_{C,i}$, and $\sigma_{C,i}$, respectively. If we denote $\langle q_{C,1},z_{C,2},\ldots,z_{C,s}\rangle$ by $I_{C}$, then the ring $T_{C}=R[x_{1},\ldots,x_{s}]/I_{C}$ is a chain ring with maximal ideal $\mathfrak{M}=\langle a,q_{C,1},z_{C,2},\ldots,z_{C,s}\rangle+I_{C}$, and the residue field $T_{C}/\mathfrak{M}\cong\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{s})$. Let us define the polynomial
@@
    If $\lvert\mathcal{C}\rvert=k$ and $\{C_{i}\}_{i=1}^{k}$ is a set of representatives $\mathcal{C}$, then the ring $\mathscr{R}$ can be decomposed as a direct sum
@@
A ring $A$ is called uniserial if all $A$-submodules of the module $A_{A}$ are linearly ordered by inclusion. A ring is called serial if it is a direct sum of uniserial rings. Any uniserial ring is local, and a serial ring is uniserial if and only if it is local, see [31, Section 1.4]. Note that any chain ring is a serial ring, thus, if $g(y)\in R[y]$ is a monic basic irreducible polynomial, then the quotient ring $R[y]/\langle g(y)\rangle$ is a serial ring. Finite commutative semi-simple rings are those rings that can be represented as finite direct sums of finite fields. The following proposition is straightforward from the previous paragraph.

### v3

@@ Abstract
In this paper, we investigate the algebraic structure for polycyclic codes over a specific class of serial rings, defined as $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ for $i\in\{1,\ldots,s\}$ is a monic square-free polynomial. We define quasi-$s$-dimensional polycyclic codes and establish an $R$-isomorphism between these codes and polycyclic codes over $\mathscr{R}$. We provide necessary and sufficient conditions for the existence of annihilator self-dual, annihilator self-orthogonal, annihilator linear complementary dual, and annihilator dual-containing polycyclic codes over this class of rings. We also establish the CSS construction for annihilator dual-preserving polycyclic codes over Frobenius rings.
@@
journal: Cryptography and Communicationsmmfootnotetext: Maryam Bajalan is supported by the Bulgarian Ministry of Education and Science, Scientific Programme “Enhancing the Research Capacity in Mathematical Sciences (PIKOM)”, No. DO1-67/05.05.2022.vvfootnotetext: Both authors were partially supported by Grant TED2021-130358B-I00 funded by MCIU/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR”. This research was partially conducted during a research stay of the second author funded by the Institute of Mathematics and Informatics at the Bulgarian Academy of Sciences.
@@ 1 Introduction
Polycyclic codes are ideals in the quotient of the polynomial ring $A[x]/\langle f(x)\rangle$, where $A$ is a ring and $f(x)$ is a polynomial in $A[x]$, see [1, 8, 9, 28]. Cyclic codes and constacyclic codes over $A$ are particular instances, when $f(x)=x^{n}-1$ and $f(x)=x-\lambda$, respectively. A ring is called uniserial if all its submodules are linearly ordered by inclusion, and it is called serial if it is a direct sum of uniserial rings. In this paper, we aim to investigate polycyclic codes in the case $A$ is the serial ring $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial for $i=1,2,\ldots,s$.
We briefly review the vast literature on polycyclic codes over some special classes of rings. Alkenani et al. [3] studied constacyclic codes over the ring $\mathbb{F}_{q}[u_{1},u_{2}]/\langle u_{1}^{2}-u_{1},u_{2}^{2}-u_{2},u_{1}u_{2}-u_{2}u_{1}\rangle$. Dinh et al. [14] investigated constacyclic codes over the ring $\mathbb{F}_{p}+u_{1}\mathbb{F}_{p}+\cdots+u_{s}\mathbb{F}_{p}$, subject to the conditions $u_{i}^{2}=u_{i}$, $u_{i}u_{j}=u_{j}u_{i}=0$, and $p$ is an odd prime. Dinh et al. [16] explored the algebraic structure of constacyclic codes over finite semi-simple rings, defined as direct sums of finite fields. Islam et al. [25] focused their study on cyclic codes over the non-chain ring $\mathbb{F}_{q}[u]/\langle u^{e}-1\rangle$. Bag et.al [6] examined skew constacyclic codes over the ring $\mathbb{F}_{q}[u,v]/\langle u^{2}-1,v^{2}-1,uv-vu\rangle$. Ali et al. [2] worked out the case cyclic codes over non-chain rings $\mathbb{F}_{q}[u_{1},u_{2},\ldots,u_{s}]/\langle u^{2}_{i}-(\alpha_{i})^{2},\quad u_{i}u_{j}-u_{j}u_{i}\rangle$, where each $\alpha_{i}$ is a non-zero element of $\mathbb{F}_{q}$. Goyal and Raka [19, 20] explored polyadic cyclic and constacyclic codes over the non-chain ring $\mathbb{F}_{q}[u,v]/\langle f(u),g(v),uv-vu\rangle$, with $f(u)$ and $g(v)$ splitting into distinct linear factors over $\mathbb{F}_{q}$. Bhardwaj et al. [10] investigated constacyclic codes over the finite non-chain ring $\mathbb{F}_{q}[u_{1},u_{2},\ldots,u_{k}]/\langle f_{i}(u_{i}),u_{i}u_{j}-u_{j}u_{i}\rangle$, where each $f_{i}(u_{i})$ is a polynomials that splits into distinct linear factors over $\mathbb{F}_{q}$. Qi [37] focused on polycyclic codes over $\mathbb{F}_{q}+u\mathbb{F}_{q}$, with $u^{2}=0$. Karthick [26] investigated polycyclic codes over the ring $R=\mathbb{F}_{q}+u\mathbb{F}_{q}+v\mathbb{F}_{q}$, subject to the conditions $u^{2}=\alpha u$, $v^{2}=v$, and $uv=vu=0$, where $\alpha$ represents a unit element of $R$.
@@
In this paper, we investigate polycyclic codes over the general ring $\mathscr{R}$. The outline of the paper is as follows. Section 2 shows the foundational structure of $\mathscr{R}$. In Section 3, we examine how a polycyclic code $\mathscr{C}$ over $\mathscr{R}$ decomposes into a direct sum $\oplus_{i=1}^{k}\mathscr{C}_{i}$, where each $\mathscr{C}_{i}$ constitutes a polycyclic code over a chain ring. Section 4 explores the concept of quasi-2-dimensional polycyclic (Q2DP) codes along with their multidimensional counterparts (QsDP), which extend the ideas of quasi-polycyclic codes and multidimensional cyclic codes [22, 23, 24, 32, 40]. Furthermore, we define the concept of $f(x)$-polycyclic-QsDP codes as a specific subclass of QsDP codes. Then, we prove that when considering polycyclic codes over $\mathscr{R}$ as $R$-submodules, they precisely coincide with $f(x)$-polycyclic-QsDP codes. Polycyclic codes are defined as the invariant $\mathscr{R}$-submodules of $\mathscr{R}^{n}$ under the operator $\tau_{f(x)}$. Their counterpart are sequential codes, which are defined as invariant subspace under the adjoint operator of $\tau_{f(x)}$. It is well-known that if a linear subspace is invariant under one operator, its Euclidean dual is invariant under the adjoint operator. Consequently, a polycyclic code $\mathscr{C}$ and its Euclidean dual $\mathscr{C}^{\perp}$ do not share the same structure. Thus, the first ones are equivalent to ideals in $\mathscr{R}[x]/\langle f(x)\rangle$, defining a polynomial and ideal structure for the second ones is challenging. Note that, if $\mathscr{C}$ is an ideal in $\mathscr{R}[x]/\langle f(x)\rangle$ generated by the polynomial $g(x)$ with the property that $f(x)=g(x)h(x)$, it cannot be assumed that $\mathscr{C}^{\perp}$ is generated by $h(x)$ because $\mathscr{C}^{\perp}$ is not necessarily a polycyclic code. This incorrect interpretation happened in the proof of Theorem 5 in [35]. In Section 5, using multivariable codes over chain rings in [29, 30], we establish a polynomial structure for $\mathscr{C}^{\perp}$ for the special case where $f(x)$ is a polynomial with coefficients in $R$. The annihilator dual is an alternative duality concept ensuring that the dual of polycyclic codes has a polycyclic structure, a feature that fails in the Euclidean dual. In Section 6, some properties of the annihilator dual of a polycyclic code over $\mathscr{R}$ are studied. Usually, it is stated in the literature that “a polycyclic code $\langle g(x)\rangle$ is annihilator self-dual if and only if $f(x)=g^{2}(x)$. As explained in Remark 5, that statement relies on the assumption that “if $b$ divides $c$ and $c$ divides $b$, then $b$ equals $c$’, without considering that the divisibility relation for rings is different, namely “if $b$ divides $c$ and $c$ divides $b$, then $\langle b\rangle$ equals $\langle c\rangle$. Thus in Section 6, we revise this property as follows, “a polycyclic code $\langle g(x)\rangle$ is annihilator self-dual if and only if $f(x)=ag^{2}(x)$ for some unit element $a$ in the ring $\mathscr{R}$. To finish this paper, we study some quantum code constructions from this type of codes. Calderbank-Shor-Steane (CSS) construction allows for constructing quantum codes from dual-preserving classical codes, see for example [34, 38]. Existing literature has applied this construction only to codes that preserve Euclidean or Hermitian duality. In Section 7, we present this construction for polycyclic codes over Frobenius rings that preserve annihilator duality.
List of symbols and abbreviations
  • $R$
    : Finite chain ring with the maximal ideal $\mathfrak{m}=\langle\gamma\rangle$ and the nilpotency index $e$.
  • $\mathbb{F}_{q}$
    : Residue field $\mathbb{F}_{q}=R/\mathfrak{m}$, where $q=p^{t}$, $p$ a prime number.
  • $\overline{\phantom{r}}$
    : Natural ring homomorphism from $R[y]$ to $\mathbb{F}_{q}[y]$.
  • $\mathscr{R}$
    : Ring $R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $t_{i}(x_{i})\in R[x_{i}]$ is a monic square-free of degree $m_{i}$.
  • $H_{i}$
    : Set of roots of $\bar{t}_{i}(x_{i})$ in an extension of $\mathbb{F}_{q}$.
  • $C(\nu)$
    : Set of the class of $\nu\in\mathcal{H}=\prod_{i=1}^{s}H_{i}$ as $\{(\nu^{q^{j}_{1}},\ldots,\nu^{q^{j}_{s}})\mid j\in\mathbb{N}\}$.
  • $\mathcal{C}$
    : Set of all classes $C(\nu)$.
  • $C$
    : Simpler notation for $C(\nu)$.
  • $p_{C,i}(x_{i})$
    : Polynomial $\text{Irr}(\nu_{i},\mathbb{F}_{q})$, where $\nu=(\nu_{1},\ldots,\nu_{s})\in\mathcal{H}$.
  • $b_{C,i}(x_{i})$
    : Polynomial $\text{Irr}(\nu_{i},\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1}))$.
  • $\widetilde{b}_{C,i}(x_{i})$
    : Polynomial $p_{C,i}(x_{i})/b_{C,i}(x_{i})$.
  • $w_{C,i}(x_{1},\ldots,x_{i})$
    : Polynomial obtained by replacing $\nu_{i}$ with $x_{i}$ in $b_{C,i}(x_{i})$.
  • $\pi_{C,i}(x_{1},\ldots,x_{i})$
    : Polynomial obtained by replacing $\nu_{i}$ by $x_{i}$ in $\widetilde{b}_{C,i}(x_{i})$.
  • $q_{C,i}$
    : Hensel lift to $R$ of $p_{C,i}$.
  • $z_{C,i}$
    : Hensel lift to $R$ of $w_{C,i}$.
  • $\sigma_{C,i}$
    : Hensel lift to $R$ of $\pi_{C,i}$.
  • $I_{C}$
    : Ideal $\langle q_{C,1},z_{C,2},\ldots,z_{C,s}\rangle$.
  • $T_{C}$
    : Chain ring $R[x_{1},\ldots,x_{s}]/I_{C}$.
  • $h_{C}$
    : Polynomial $h_{C}(x_{1},x_{2},\ldots,x_{s})=\prod_{i=1}^{s}t_{i}(x_{i})q_{C,i}(x_{i})\prod_{i=1}^{s}\sigma_{q_{C,i}}(x_{2},\ldots,x_{i})$.
  • $e_{C}$
    : Orthogonal idempotent such that $\mathscr{R}e_{C}\cong\langle h_{C}+I\rangle$.
  • $\mathscr{R}_{i}$
    : Chain ring isomorphic to $\mathscr{R}e_{C_{i}}$, where $\{C_{i}\}_{i=1}^{k}$ is a set of representatives of $\mathcal{C}$.
  • $\mathscr{C}$
    : Linear code of length $n$ over $\mathscr{R}$.
  • $f(x)$
    : Polynomial $x^{n}-(f_{n-1}x^{n-1}+\cdots+f_{1}x+f_{0})$ in $\mathscr{R}[x]$.
  • $\tau_{f(x)}$
    : $f(x)$-polycyclic shift.
  • $\boldsymbol{c}$
    : Element $(c_{0},c_{1},\ldots,c_{n-1})\in\mathscr{R}^{n}$.
  • $c_{i,j}$
    : Components of $c_{j}$ ($10\leqslant j\leqslant n-$ 1) in $c_{j}=c_{1,j}e_{C_{1}}+c_{2,j}e_{C_{2}}+\cdots+c_{k,j}e_{C_{k}}\in\mathscr{R}$.
  • $\boldsymbol{c}_{i}$
    : Components of $\boldsymbol{c}$ in $\boldsymbol{c}=\boldsymbol{c}_{1}e_{C_{1}}+\boldsymbol{c}_{2}e_{C_{2}}+\cdots+\boldsymbol{c}_{k}e_{C_{k}}$.
  • $\mathscr{C}_{i}$
    : Set $\{\boldsymbol{c}_{i}\in\mathscr{R}^{n}\mid\boldsymbol{c}_{1}e_{C_{1}}+\boldsymbol{c}_{2}e_{C_{2}}+\cdots+\boldsymbol{c}_{k}e_{C_{k}}\in\mathscr{C}\}$.
  • $f_{i}(x)$
    : Polynomial $x^{n}-(f_{i,n-1}x^{n-1}+\cdots+f_{i,1}x+f_{i,0})$, where $f_{i,j}$ is component of $f_{j}$.
  • $\operatorname{rank}_{\mathscr{R}}(\mathscr{C})$
    : Minimum number of generators for spanning $\mathscr{C}$ as $\mathscr{R}$-linear combinations.
  • $\mathscr{R}_{f}$
    : Quotient ring $\mathscr{R}[x]/\langle f(x)\rangle$.
  • $\phi$
    : Map that associates each vector in $\mathscr{R}$ with its corresponding polynomial in $\mathscr{R}_{f}$.
  • $\mathbf{x}^{\alpha}$
    : Monomial $x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\ldots x_{s}^{\alpha_{s}}$, where $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{s})\in(\mathbb{N}\cup\{0\})^{s}$.
  • $\succ$
    : Lexicographic ordering on the monomials.
  • $\Phi$
    : $R$-isomorphism from $\mathscr{R}$ to $R^{m}$, where $m=m_{1}m_{2}\ldots m_{s}$.
  • $QsDP$
    : Quasi-s-dimensional polycyclic codes.
  • $C_{f}$
    : Companion matrix of the polynomial $f$.
  • $I_{i\times i}$
    : Identity matrix.
  • $\otimes$
    : Tensor product of matrices.
  • $\nu_{f(x)}$
    : $f(x)$-sequential shift.
  • $\mathscr{C}^{\perp}$
    : Euclidean dual of $\mathscr{C}$.
  • $\mathscr{C}^{\circ}$
    : Annihilator dual of $\mathscr{C}$.
  • $LCD$
    : Linear complementary dual code.
  • $\chi$
    : Character in $\operatorname{Hom}(R;\mathbb{C}^{*})$.
  • $\operatorname{Ann}(\mathscr{C})$
    : Annihilator of the subset $\mathscr{C}$ of $\mathscr{R}_{f}$.
  • $\operatorname{swt}$
    : Symplectic weight.
  • $\operatorname{wt}$
    : Hamming weight.
@@
Let $R$ be a finite commutative chain ring, and let $\gamma\in R$ be a fixed generator of the maximal ideal $\mathfrak{m}$ with nilpotency index $e$. Consider $\mathbb{F}_{q}=R/\mathfrak{m}$ as the residue field, where $q=p^{t}$, $p$ a prime number. Let ¯$\overline{\phantom{x}}:R[y]\to\mathbb{F}_{q}[y]$ represent the natural ring homomorphism, mapping $r\in R$ to $r+\mathfrak{m}$ and the variable $y$ to itself. A polynomial $p(y)=\sum_{i=0}^{n-1}p_{i}y^{i}\in R[y]$ is called square-free (basic irreducible) if $\bar{p}(y)$ is square-free (basic irreducible) in $\mathbb{F}_{q}[y]$. Additionally, it is called regular if it is not a zero divisor in $R[y]$. In [15, Proposition 2.3] it is stated that $p(y)$ is regular if and only if $p_{i}$ is a unit element for some $i$. Hence, monic polynomials are regular. If $g(y)\in R[y]$ is a regular basic irreducible polynomial, then the quotient ring $R[y]/\langle g(y)\rangle$ forms a chain ring, see [15, Lemma 3.1]. The monic square-free polynomial $g(y)$ factors uniquely as a product of monic basic irreducible pairwise coprime polynomials, given by $g(y)=g_{1}(y)g_{2}(y)\cdots g_{r}(y)$ [33, Theorem XIII.4]. Hence, by the Chinese Remainder Theorem, the quotient ring $R[y]/\langle g(y)\rangle$ can be decomposed as a direct sum of chain rings
@@
The algebraic properties of the algebra $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, assuming that each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial of degree $m_{i}$, was explored in [29]. Let $H_{i}$ denote the set of roots of $\bar{t}_{i}(x_{i})$ in a suitable extension of $\mathbb{F}_{q}$. For each $\nu\in\mathcal{H}=\prod_{i=1}^{s}H_{i}$, let $C(\nu)$ denote the set of the class of $\nu$ as $\{(\nu^{q^{j}_{1}},\ldots,\nu^{q^{j}_{s}})\mid j\in\mathbb{N}\}$. The set of all these classes, denoted by $\mathcal{C}=\mathcal{C}(t_{1},\ldots,t_{s})$, forms a partition of $\mathcal{H}$.
The size of each class is given by $|C(\nu)|=\text{lcm}(d_{1},\ldots,d_{s})=[\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{s}):\mathbb{F}_{q}]$, where $d_{i}$ is the degree of the irreducible polynomial of $\nu_{i}$ over $\mathbb{F}_{q}$. Let $p_{C,i}(x_{i})$ denote the polynomial $\text{Irr}(\nu_{i},\mathbb{F}_{q})$, where $\nu=(\nu_{1},\ldots,\nu_{s})\in\mathcal{H}$ and $C=C(\nu)$ is a class of $\nu$. For each $i=2,\ldots,s$, let $b_{C,i}(x_{i})=\text{Irr}(\nu_{i},\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1}))\in\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{i-1})[x_{i}]$ and $\widetilde{b}_{C,i}(x_{i})=p_{C,i}(x_{i})/b_{C,i}(x_{i})$. The polynomials $b_{C,i}(x_{i})$ and $\widetilde{b}_{C,i}(x_{i})$ are coprime polynomials. One can define the multivariable polynomials $w_{C,i}(x_{1},\ldots,x_{i})$ and $\pi_{C,i}(x_{1},\ldots,x_{i})$ obtained by substituting $\nu_{i}$ by $x_{i}$ in polynomials $b_{C,i}(x_{i})$ and $\widetilde{b}_{C,i}(x_{i})$, respectively and denote the Hensel lifts to $R$ of the polynomials $p_{C,i}$, $w_{C,i}$, and $\pi_{C,i}$ by $q_{C,i}$, $z_{C,i}$, and $\sigma_{C,i}$, respectively. If we denote $\langle q_{C,1},z_{C,2},\ldots,z_{C,s}\rangle$ by $I_{C}$, then the ring $T_{C}=R[x_{1},\ldots,x_{s}]/I_{C}$ is a chain ring with maximal ideal $\mathfrak{M}=\langle a,q_{C,1},z_{C,2},\ldots,z_{C,s}\rangle+I_{C}$, and the residue field $T_{C}/\mathfrak{M}\cong\mathbb{F}_{q}(\nu_{1},\ldots,\nu_{s})$. Let us define the polynomial
@@
    If $\lvert\mathcal{C}\rvert=k$ and $\{C_{i}\}_{i=1}^{k}$ is a set of representatives of $\mathcal{C}$, then the ring $\mathscr{R}$ can be decomposed as a direct sum
@@
A ring $A$ is called uniserial if all $A$-submodules of the module $A_{A}$ are linearly ordered by inclusion. A ring is called serial if it is a direct sum of uniserial rings. Any uniserial ring is local, and a serial ring is uniserial if and only if it is local, see [36, Section 1.4]. Note that any chain ring is a serial ring, thus, if $g(y)\in R[y]$ is a monic basic irreducible polynomial, then the quotient ring $R[y]/\langle g(y)\rangle$ is a serial ring. Finite commutative semi-simple rings are those rings that can be represented as finite direct sums of finite fields. The following proposition is straightforward from the previous paragraph.